a sum of money amounts to rs 13230 in 1 year and rs 13891.50 in 1.5 year at compound intrest compounded semi anually. Find the sum and the rate of intrest per anum.

Grade:11

1 Answers

Arun
6 years ago
let the sum  = p
rate of interest = r %
A1 = 13230
A2 = 13891.50
rate of interest in half yearly will be  = r/2 % (conversion period)
as the amount is invested for 1 year and 1.5 years
 so conversion are 2 and 3
 
A = P(1 + r/100)^n
13230 = P(1 + r/200)^2    …..........(i)
13891.5 = P(1 + r/200)^3 …...........(ii)
divide (ii) by (i)
 1 + r/200 = 1.05
r/200 = .05
r = 10 %
put the value of r in  (i) or (ii) – 
P = 12000 rupees